# 2. In a given country, only two good on the production possibility frontier: s are produced: potatoes or airplanes. Here are the points Maximum Annual Output Options Potatoes Airplanes 1,300 1,100 900 700 500 350 450 525 575 630 (i) Draw a production possibility frontier with the quantity of potatoes on the horizontal (Gi) What is the opportunity cost per unit of expending production from 450 airplanes to 525 (ii)at is the opportunity cost per unit of increasing the annual output of potatoes from 700 axis airplanes? to 900? (iv) Why is the production possibility graph negatively sloped? Provide a clear example. (v) Why is it bowed out (concave down)? Provide a clear example. (vi)is the opportunity cost constant? Why?
Answer #1

i) The image of this answer is attached here.6-S 535 450 350 O 5eo lo 900 lloo 1300 a pota es

ii) The opportunity cost per unit of expanding the production from 450 units of aeroplanes to 525 units is = number of units of potatoes lost / number of units of airplanes gained

= 200/ 75 = 2.67. The number of units of potatoes lost = 900 -700 on moving from point C to point D.

iii) The opportunity cost per unit of increasing the annual output of potatoes from 700 to 900 units is = number of units of aeroplanes lost/ number of units of potatoes gained

= 75/200 = 0.375. The number of potatoes gained 75 = 525- 450 on moving from point D to point C .

iv) The production possibility graph shows the combination of goods and services in the economy that has been produced efficiently in the economy at that point of time, using the available input. Here, for example to produce more units of potatoes on the production possibility curve, the production of airplanes must decrease. If an economy is producing the goods efficiently then it is not possible to produce one good without decreasing the production of another good. The negative slope of the production possibility curve indicates the opportunity cost of the goods. Production of one good cannnot be increased without decreasing the production of another good as goods are efficiently produced on the production possibility curve. Suppose, here from the given graph, we are at point C producing 900 units of potato and 450 units of aeroplane. If we want to move from point C to point B, then we increase the production of potatoes to 1100 from 900 but at the cost of aeroplanes. The production of aeroplane decreases by 100 units from 450 units to 350 units. Therefore, the opportunity cost of producing 200more units of potatoes at point C is the reduction of aeroplane by 100 units. Hence, the curve is negatively sloping as to produce one extra unit of any of the two goods the other good has to be reduced.

(v) The production possibility curve is concave in origin to show increased opportunity cost with increased output of the good. The slope of the curve at any point in time is called the marginal rate of transformation (MRT). The slope also defines the rate at which the resources can be redirected or r allocated to the production of another good. It is also commonly kno n and the opportunity cost. The opportunity cost is increasing in this case of concave production possibility curve. The shape of the ppf is concave to show the increased opportunity cost of the goods. Hence, the MRT increases in absolute value from the top of the curve till the bottom of the curve. From this graph given here, we can see for every 200 units reduction of potatoes the percentage of increase in the output of aeroplanes decreases. Hence,the ppf is concave in shape.

(v) In the case of a concave production possibility curve, the opportunity cost is increasing, in case of a convex one the opportunity cost is decreasing and that of a straight one the opportunity cost is constant. In case of constant opportunity cost, for every increase in the production of good 'X' the decrease in the production of good 'Y' will be the same.